OpenStudy (fibonaccichick666):

Can you check my answer? Calculate \(e^{tA}\) with A={3,-2;2,-2} by the method of your choice.

OpenStudy (fibonaccichick666):

I got \[e^{tA}=\left[\begin{matrix}{\frac{-1}{3}e^{2t}+\frac{4}{3}e^{-t}} & {e^{2t}-e^{-t}} \\ {-e^{2t}+e^{-t}} & {\frac{4}{3}e^{2t}-\frac{1}{3}e^{-t}}\end{matrix}\right]\]

OpenStudy (fibonaccichick666):

@ganeshie8 , do you mind just a quick check?

OpenStudy (anonymous):

its my fav subject :3

OpenStudy (fibonaccichick666):

good to know

ganeshie8 (ganeshie8):

matrix exponentials! @dan815 still rememeber these from LA ?

OpenStudy (fibonaccichick666):

haha

ganeshie8 (ganeshie8):

are there any initial conditions ?

OpenStudy (fibonaccichick666):

nope that''s it. I'm scanning my work now

OpenStudy (anonymous):

ok , i''ll check for u :D

OpenStudy (fibonaccichick666):

here is my work

ganeshie8 (ganeshie8):

wolfram agrees with your answer http://gyazo.com/5cabd46e2faec6d3336bf6fc4bb405ef

OpenStudy (fibonaccichick666):

But I don't have any of the c's

ganeshie8 (ganeshie8):

*more or less it matches upto a constant, but i found this solution by solving the system without initial conditions

OpenStudy (fibonaccichick666):

but there are no initial conditions

ganeshie8 (ganeshie8):

let me pull up my LA notes

OpenStudy (fibonaccichick666):

But my notes don't have a c either

OpenStudy (anonymous):

oh u used Laplace inverse ,using psi function is much easy

OpenStudy (fibonaccichick666):

we didn't learn that

OpenStudy (fibonaccichick666):

(I don't think so at least)

OpenStudy (anonymous):

kk , its shortcut way for ur solution :) since u dint take yet u wont understand it until u've prove things first xD so leave it, ur solution looks good to me :|

OpenStudy (fibonaccichick666):

awesome, thank you. Out of curiosity, what is the psi way?

OpenStudy (anonymous):

its a function been used as solution to solve some cases of ODE , homogeneous and non

OpenStudy (anonymous):

to get e^tA , we only need to find eigen values and vectores with some trick from psi

OpenStudy (fibonaccichick666):

darn, that would have been easier